2. The average length of time for students to


have their subjects controlled is 40 minutes. A


new controlling procedure using modern


computing machines is being tried. If a random


sample of 15 students has an average


controlling time of 25 minutes with a standard


deviation of 12.9 minutes under the new


system, test the hypothesis that the average


length of time to control student’s subjects is


less than 40 minutes. Use a level of


significance of 0.10 and assume the


population of controlling times to be normally


distributed.

1
Expert's answer
2022-05-17T09:48:12-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge 40"

"H_a:\mu<40"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.1," "df=n-1=14" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.34503."

The rejection region for this left-tailed test is "R = \{t: t < -1.34503\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{25-40}{12.9/\sqrt{15}}\approx-4.50347"

Since it is observed that "t = -4.50347 <-1.34503= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed "df=14" degrees of freedom, "t=-4.50347" is "p = 0.000248," and since "p=0.000248<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 40, at the "\alpha = 0.10" significance level.